40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 26 01:01:34 2010, Zach, Laser, GYRO, gyro back to gyro configuration 2010-08-25_13.19.11.jpg2010-08-25_13.20.31.jpg
    Reply  Fri Aug 27 11:51:41 2010, Zach, Laser, GYRO, gyro back to gyro configuration 
Message ID: 986     Entry time: Thu Aug 26 01:01:34 2010     Reply to this: 997
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: gyro back to gyro configuration 

 I put the gyro back into the normal configuration today. It is now locking in both directions again. The bandwidth seems slightly improved, as I didn't have to be quite as careful when working on the table to keep it from dropping, but it is far from good, and the 200kHz+ oscillation still shows up with enough gain. I had set up the PLL but was having problems locking it when I had to leave.

A couple changes:

  • I kept the high-speed PDA10A that I used for the linear cavity in place as the CCW REFL PD, instead of the old PDA36A. We eventually want to have the same two PDs for both REFLs, but this doesn't really matter at the moment, and we might as well maximize our optical gain in the main locking loop while using the design modulation frequency of 33 MHz. The CW loop is still a 17-MHz diode, but this loop seems to like the lowest gain setting anyway, and we may end up wanting to reduce lower gain limit of the CW PDH box.
  • The BS1 in the TRANS readout was a 45P, but we are using S, so I changed it to a 45S.

I took some time to label nearly all of the cables we work with at the controls interface. The REFL PD signals, the CW & CCW error signals, the PZT and AOM VCO control signals, the TRANS DC signal, the SWEEP drive from the function generator, and two cables going to dedicated DAQ channels have all been labeled. This way we don't have to tug on cables or follow them across the table to figure out what is what. We should be prudent that we don't change them around without relabeling them appropriately

2010-08-25_13.19.11.jpg2010-08-25_13.20.31.jpg

Tomorrow morning I'm going to figure out what's wrong with the PLL, then lock it and take a spectrum to see if things have gotten better with the new PD, new modulation frequency, and fresh realignment. Then I plan to take a bunch of transfer functions to try and diagnose our loop problem. Specifically:

  1. The closed-loop TF of each direction
  2. The OLTF from the SWEEP input of each PDH box to its output

(2) is necessary to obtain the OLTF of the loops from (1), as we will be using the SWEEP input for the excitation, and the TF from SWEEP to OUT is not the same as that from IN to OUT.

ELOG V3.1.3-